OpenStudy (anonymous):

Enter the quadratic equation in standard form in the box below. x^2-10x+16=-5

OpenStudy (anonymous):

Do you know the standard form for a quadratic equation?

OpenStudy (anonymous):

ax2 + bx + c = 0

OpenStudy (anonymous):

So just express the given equation in the above format.

OpenStudy (shubhamsrg):

this is amazing!! a cardboard seems smarter than you!! i've told you these pathetic ques twice i guess..

OpenStudy (anonymous):

Please calm down, we must try and make him understand.

OpenStudy (shubhamsrg):

do your damn homework yourself!!

OpenStudy (shubhamsrg):

sir this guy is amazing,,you should see his previous ques..and then compare with this one..

OpenStudy (anonymous):

You don't need to call me sir, I'm just in ninth grade.

OpenStudy (shubhamsrg):

see this betta..he with his female version :P http://openstudy.com/study?signup#/updates/4fad2890e4b059b524f8ae73

OpenStudy (shubhamsrg):

ohh..lol..alright..

OpenStudy (anonymous):

OK, @marcoduuuh , do you know how to make the RHS of the equation 0?

OpenStudy (shubhamsrg):

buddy you're just wasting time,,anyways,,am outa here..

OpenStudy (anonymous):

So do you know?

OpenStudy (anonymous):

I'll do my best to help you understand this.

OpenStudy (anonymous):

x^2-10x+21=0 (x-7)(x-3)=0 x=7 & x=3 ?

OpenStudy (anonymous):

Do you want to find the value of x? I'm guessing that you're confused how there are two values of x.

OpenStudy (anonymous):

Yeaaah pretty much.

OpenStudy (anonymous):

OK, now this is a quadratic equation, which means the degree is 2.You get that?

OpenStudy (anonymous):

Oh, wait, I think I got it, is it this? 9-30+16=-5?

OpenStudy (anonymous):

No, you don't need to confuse yourself.

OpenStudy (anonymous):

I'm explaining why there are two values of x. You understand the part about the degree.

OpenStudy (anonymous):

Yeaah sure!

OpenStudy (anonymous):

OK, so basically, the concept is that the degree of the equation denotes the number of solutions it has. You interpret it like this, a quadratic equation will have two values, an equation with degree 3 will have 3 solutions and so on.

OpenStudy (anonymous):

Okaaay! Thank you @pratu043 I think I got this!

OpenStudy (anonymous):

I suppose that clears up your doubts.

OpenStudy (anonymous):

Indeed! Thank you!

OpenStudy (anonymous):

x^2-10x+21=0 right?

OpenStudy (anonymous):

That's right. Sorry for the delay and welcome!!